OpenStudy (anonymous):

Jose is paying for a $10 meal using bills in his wallet. He has five $1 bills and four $5 bills. If he selects two bills at random, one at a time from his wallet, without replacement, what is the probability that he will choose two $5 bills to pay for the meal? Show your work.

OpenStudy (anonymous):

\begin{array}l\color{#FF0000}{\text{RAINBOWRAINBOWRAINBOWRAINBOW}}\end{array}\begin{array}l\color{#FF7F00}{\text{RAINBOWRAINBOWRAINBOWRAINBOW}}\end{array}\begin{array}l\color{#FFFF00}{\text{RAINBOWRAINBOWRAINBOWRAINBOW}}\end{array}\begin{array}l\color{#00FF00}{\text{RAINBOWRAINBOWRAINBOWRAINBOW}}\end{array}\begin{array}l\color{#0000FF}{\text{RAINBOWRAINBOWRAINBOWRAINBOW}}\end{array}\begin{array}l\color{#6600FF}{\text{RAINBOWRAINBOWRAINBOWRAINBOW}}\end{array}\begin{array}l\color{#8B00FF}{\text{RAINBOWRAINBOWRAINBOWRAINBOW}}\end{array}

OpenStudy (anonymous):

wow

OpenStudy (aravindg):

aa

OpenStudy (aravindg):

wow